### v2

# Personalisation via Dynamic Policy Fusion

Abstract
Deep reinforcement learning (RL) policies, although optimal in terms of task rewards, may not align with the personal preferences of human users. To ensure this alignment, a naïve solution would be to retrain the agent using a reward function that encodes the user’s specific preferences. However, such a reward function is typically not readily available, and as such, retraining the agent from scratch can be prohibitively expensive. We propose a more practical approach - to adapt the already trained policy to user-specific needs with the help of human feedback. To this end, we infer the user’s intent through trajectory-level feedback and combine it with the trained task policy via a theoretically grounded dynamic policy fusion approach. As our approach collects human feedback on the very same trajectories used to learn the task policy, it does not require any additional interactions with the environment, making it a zero-shot approach. We empirically demonstrate in a number of environments that our proposed dynamic policy fusion approach consistently achieves the intended task while simultaneously adhering to user-specific needs.
Keywords:
Reinforcement Learning, Personalisation, Dynamic Policy Fusion
1 Introduction
Reinforcement learning (RL) has demonstrated its effectiveness in various real-world scenarios, including computer games [16], inventory management, autonomous driving, robotics [8], healthcare [30], recommendation systems [25] etc. In RL, the learning agent typically learns to maximise its reward by learning an optimal task policy through interactions with the environment. However, in certain scenarios, a user may desire a policy that subtly deviates from this optimal policy to accommodate their personal preference or style. For example, in a navigation setting, users may want to avoid toll roads, or they may prefer to drive along a scenic coastal route, even if this translates to slightly longer travel times compared to the route determined to be the optimal one by an RL agent trained to minimise travel time.
A simple solution to accommodate such preferences in a policy would be to retrain an RL agent from scratch, using a reward function that takes users’ preferences into account. However, such a reward function could be challenging to obtain. In addition, the poor sample efficiency of RL could render this approach infeasible, as it would require the agent to interact with the environment again, which could be expensive in terms of time, energy, etc., We posit that a more practical approach would be to adapt the originally trained policy by personalising it based on a user’s specific preferences on the trajectories already collected during training. This approach to personalisation eliminates the need for additional agent-environment interactions, thereby making it a practically feasible option for adapting already trained policies.
In this work, we present a policy fusion-based approach to learn such a Personalised policy that respects human preferences while also satisfying the task objective. To infer human preferences, the trajectories used to initially learn the task (we refer to the corresponding policy as the Task-specific policy) are numerically scored to reflect the degree to which they align with the human’s preferred behaviour. These trajectory-score pairs are fed to an LSTM [10] to construct an Intent-specific policy. We choose an LSTM-based approach as it is well-suited to model temporal patterns within the trajectories. Once the user’s intent is inferred, the key idea is then to fuse the task-specific policy with this intent-specific policy, such that the resulting personalised policy completes the task at hand while simultaneously respecting user-specific needs. We design the described policy fusion method to theoretically ensure that the divergence of the personalised policy with respect to the task-specific policy remains bounded implying that the personalised policy upon fusion cannot change arbitrarily. Since the inference of the user’s intent (learning the intent-specific policy) is done by reusing the same trajectories used for training the task policy/task-specific policy, our approach obviates the need for any additional environment interactions, making it a zero-shot scheme.
Despite the mentioned advantages of the described static policy fusion-based approach, we show that this policy fusion approach is insufficient, as it does not account for whether one policy is over/under-represented, which can lead to undesirable effects. For example, in a navigation scenario, the preference for visiting a particular state may cause the agent to repeatedly visit that state, thereby ignoring the primary navigation objective. In other words, static policy fusion may cause one of the policies to dominate the other, leading to unwanted agent behaviours. We address this problem through a novel dynamic policy fusion approach that automatically regulates the dominance of the intent-specific policy by regulating the associated temperature parameter of the Boltzmann distribution [23] used to represent the policy. This helps control the relative contribution of the intent-specific policy in the personalised policy, thereby allowing the agent to complete the task while respecting user-specific needs/preferences.
We note that the problem of respecting additional user-specific objectives could also be formulated as a multi-objective reinforcement learning [22, 9, 15] problem, although such approaches are applicable only when rewards corresponding to each objective are available during learning. As such, the present work aims to deal with how one can adapt a trained policy after it has learned an initial task. Overall, the contributions of this work can be summarised as follows:
  • A technique to achieve zero-shot personalisation by adapting an already trained task policy without additional environmental interactions.
  • Theoretical analysis establishing the boundedness of the personalised policy with respect to the task policy.
  • Demonstration of the failure of static policy fusion methods, motivating the need for dynamic fusion techniques.
  • A modulation mechanism for dynamic policy fusion to balance task performance with user preference.
2 Related Work
Learning from human feedback is of particular interest in the RL community as it leverages human knowledge during the learning process, offering several benefits. Firstly, it improves the efficiency of the system in terms of sample requirements as well as overall performance  [6]. Secondly, leveraging human feedback has been shown to enable RL agents to solve complex tasks that are otherwise challenging to manually specify through conventional reward functions [12, 5]. Consequently, there exist a number of approaches aimed at leveraging human feedback for learning.
A number of works in interactive reinforcement learning [28, 14] examine how agents can learn from state-action level human feedback. However, such methods are limited, as providing state-action-level feedback is non-intuitive and cognitively taxing. Preference-based reinforcement learning [5, 12], uses trajectory-level human preference feedback, using which a corresponding reward function is learned. However, following the learning of the reward function, this approach may require the agent to interact with the environment again to learn the corresponding policy. As our present work focuses on developing a zero-shot approach, to estimate human intent, we instead use RUDDER [1], an LSTM-based approach for credit assignment by directly approximating the Q-values from agent trajectories. These Q-values are then translated into the required intent-specific policy. There are more recent work addressing credit assignment such as [19, 31], but we resort to RUDDER for our work due to it’s simplicity.
In the context of policy fusion, a closely related work by Sestini et al. [24] combines policies with different fusion methods. However, the described policy fusion methods are static and may result in the dominance of one of the constituent policies. Policy fusion has been explored previously in Haarnoja et al. [7] and Hunt et al. [11]. Such works study the concept of learning different policies independently using soft Q-functions, each with its own reward function. These individual policies are later combined, leading to the emergence of new behaviours in robotic manipulations. Our work views policy fusion through a lens of personalisation, requiring the fusion to satisfy certain constraints. In addition, unlike existing policy fusion methods, our dynamic policy fusion approach prevents the over-dominance of the component policies.
The overarching aim of our work is to personalise existing task policies using trajectory level human feedback without the need for additional environment interactions. The goal of personalisation could possibly be viewed from a multi-objective reinforcement learning (MORL) [9] or constrained reinforcement learning [27] perspective. Multi-objective reinforcement learning (MORL) embodies personalisation by determining which objectives take precedence, and addressing conflicting objectives [2] through Pareto-optimal solutions. In Mo et al. [17], a personalised dialogue system is developed using transfer learning, adapting common dialogue knowledge from a source domain to a target user, whereas Bodas et al. [4] aims to personalise non-player characters (NPCs) to human skill levels to enhance player engagement. This approach designs a composite scalar reward function that implicitly matches the NPC’s skill level to that of the human player. However, it necessitates additional environment interaction and NPC training with the engineered reward function. Unlike these works, our approach does not require any additional interaction or specifically crafting the reward function. We believe these benefits make the present work particularly relevant for real-world scenarios.
3 Preliminaries and Background
We briefly survey some of the related backgrounds that form the basis for our work. We refer to the policy trained to solve the task as the task-specific policy $(\pi_{\phi})$ and the policy that captures human intent as the intent-specific policy$(\pi_{\psi})$.
3.1 Reinforcement Learning
The policies considered in this work are learned by solving tasks represented as individual MDPs [21] $\mathcal{M}=(\mathcal{S},\mathcal{A},P,R,\gamma)$ where $\mathcal{S}$, and $\mathcal{A}$ are the state and action spaces. $P$ is the transition function that captures the transition dynamics of the environment. $R$ is the reward function and $\gamma$ is the discount factor. In each task, $R$ can vary, while the remaining components stay the same. At timestep $t$, the agent in state $s_{t}\in\mathcal{S}$ takes an action $a_{t}\in\mathcal{A}$ and obtains a reward $R(s_{t+1},s_{t},a_{t})$ and moves to state $s_{t+1}$ according to the transition function $P(s_{t+1}|s_{t},a_{t})$. A policy $\pi(a_{t}|s_{t})$ outputs the probability of taking an action $a_{t}$ from a given state $s_{t}$. The episodic discounted return is $G_{t}=\sum_{t=0}^{T}\gamma^{t}r_{t}$, where $\gamma$ specifies how much the future reward is discounted and $T$ is the total number of timesteps. The agent’s objective is to maximise the future expected reward $\mathbb{E}[G_{t}]$ by learning a Q-function $Q(s_{t},a_{t})$, which estimates the expected cumulative reward.
3.2 RUDDER
RUDDER [1] addresses the problem of credit assignment and learning from sparse rewards. In RUDDER, an LSTM network analyses the entire trajectory data with a score received at the end to estimate Q-values for individual state-action pairs within the trajectory. These Q-values represent the expected future reward for taking an action in a given state. RUDDER also showed a densified reward can be formulated from the human feedback score as follows,
$E[r_{t+1}|s_{t},a_{t},s_{t-1},a_{t-1}]=Q^{\prime}(s_{t},a_{t})-Q^{\prime}(s_{t-1},a_{t-1}),$ (1)
where, $Q^{\prime}(s_{t},a_{t})$ represents the LSTM-generated Q-values corresponding to the state-action pair at the specific time step $t$.
We use the RUDDER framework to learn the Q-values corresponding to human scores and later convert the Q-values to policies as described in Sections 4.1 and 4.2. The densified reward as in (1) is used to modulate the influence of the policy as discussed in Section 4.5. For more details of the training of LSTM, please refer to Section 0.B of the supplementary.
4 Methodology
We construct our proposed zero-shot approach for personalisation by first inferring the intent-specific policy via trajectory-level human feedback using RUDDER, followed by a dynamic policy fusion mechanism that automatically maintains a balance between the inferred policy and a trained task-specific policy.
We assume that the trajectory data used to train task-specific policy is accessible, and reusing this data allows us to perform zero-shot personalisation i.e., without collecting new environment interactions. A subset of this data is sampled with personalised human feedback scores with more desirable trajectories assigned higher scores. We then infer the human intent (intent-specific policy) using RUDDER (Section 3.2) and dynamically fuse it with the task-specific policy. The overview of our method is illustrated in Figure 1. The subsequent sections provide in-depth details of the components of our personalisation approach.
Figure 1: Summary of our method. Trajectory-level training data used to train the task policy is labelled by a human user and an LSTM model is employed to identify the intent of the human. The output from the LSTM is converted to an intent-specific policy which is then dynamically fused by modulating the temperature parameter of the intent-specific policy (Discussed in Section 4.5) with the task-specific policy to obtain the personalised policy. The blue shaded block is the task policy training and the green dotted block is the dynamic policy fusion.
4.1 Learning Human Intent using LSTM
To learn the intent-specific policy, as previously described, we leverage the LSTM-based approach of RUDDER. We train this LSTM using human trajectory-level feedback on the same training data used to learn the task-specific policy. Hence, no additional interaction data is required. To get the human feedback, we simulated the human. The feedback score is the number of times the human preference is met within a trajectory (Refer to supplementary Section 0.A for more details).
At each time step, the state-action vector is fed into the LSTM units. In the case of image inputs, we pre-train a Variational Autoencoder (VAE) to reduce the dimensionality of the state. We use FiLM [20] to modulate the state vector with the action vector. In [20], a FiLM network conditions the feature map of the neural network depending on another input signal. Here, we condition the state feature with the action vector and train the LSTM using the human feedback data as training labels. With this training setup, (More details in supplementary Section 0.B and 0.C) the model approximates the Q-value, which is then converted to the intent-specific policy as described in the next section.
4.2 Policy Construction
We use a DQN parameterised by $\phi$ to learn the task and it produces the Q-values ($Q$) for each state-action pair. Similarly, the LSTM parameterised by $\psi$ also produces the Q-values ($Q^{\prime}$) corresponding to the human intent. We choose the Boltzmann distribution to convert the Q-values into a policy. Therefore $\pi_{\phi}$ is the task-specific policy and $\pi_{\psi}$ is the intent-specific policy. This choice is motivated by the fact that if the intersection of the support of two policies is an empty set, their product would result in a random policy. By contrast, the Boltzmann distribution assigns at least a small probability to all actions, ensuring that the support of the distribution covers the entire action space.
$\pi_{\phi}(a|s_{t})=\frac{\exp{(\frac{Q(s_{t},a)}{T_{\phi}})}}{\sum_{a\in\mathcal{A}}\exp{(\frac{Q(s_{t},a)}{T_{\phi}})}},$ (2)
$\pi_{\psi}(a|s_{t})=\frac{\exp{(\frac{Q^{\prime}(s_{t},a))}{T_{\psi}})}}{\sum_{a\in\mathcal{A}}\exp{(\frac{Q^{\prime}(s_{t},a))}{T_{\psi}})}},$ (3)
where $T_{\phi}$ is the temperature corresponding to task-specific policy and $T_{\psi}$ is the temperature corresponding to intent-specific policy. With policies constructed from Q-values, these constituent policies are fused together to produce the personalised policy
4.3 Policy Fusion
Policy fusion is a process of combining two or more policies to produce a new policy. Sestini et al. [24] discussed several approaches through which policy fusion can be performed. We also aim to obtain a personalised policy by fusing the intent-specific and task-specific policies. However, policy fusion in the context of personalisation should satisfy two constraints. Firstly, fused policy should be identical to the task-specific policy if the task-specific and intent-specific policies are identical to each other - a property that we refer to as the invariability constraint. This constraint is crucial because it ensures that when the objectives of the task and the human are aligned, the personalised policy remains consistent and does not deviate during the fusion of identical constituent policies. Correspondingly, we define invariant policies as follows:
Definition 1 (Invariant policies)
Two policies $\pi_{1}(a|s)$ and $\pi_{2}(a|s)$ are said to be invariant policies if $KL\left(\pi_{1}(a|s)\middle\|\pi_{2}(a|s)\right)=0\ \forall\ s\in\mathcal{S}$.
The second constraint is that the fused policy should act on the common support of the policies being fused. To accommodate the aforementioned constraints, we construct the personalised policy through a new fusion method as follows:
$\pi_{f}(a|s)=\tfrac{1}{Z}\sqrt{\pi_{\phi}(a|s_{t})\times\pi_{\psi}(a|s_{t})},$ (4)
where Z is the normalising factor. Since $\pi_{\phi}$ and $\pi_{\psi}$ are independent, their joint probability can be expressed as the product of $\pi_{\phi}$ and $\pi_{\psi}$. This ensures that the fused policy acts on the common support. Moreover, this fusion method ensures both the task-specific policy and the fused policy are invariant policies if the task-specific policy and the intent-specific policy are identical, thereby satisfying the invariability constraint.
Lemma 1
If $\pi_{\phi}(a|s)=\pi_{\psi}(a|s)$, then $\pi_{\phi}(a|s)$ and $\pi_{f}(a|s)$ are invariant policies.
Proof
If $\pi_{\phi}(a|s)=\pi_{\psi}(a|s)$ then $KL\left(\pi_{\phi}(a|s)\middle\|\pi_{f}(a|s)\right)=KL\left(\pi_{\phi}(a|s)\middle\|\pi_{\phi}(a|s)\right)=0$
The above lemma captures the invariability of the proposed fused method when $\pi_{\phi}(a|s)=\pi_{\psi}(a|s)$. However, if the deviation between the Q-values of the intent-specific policy and the task-specific policy is bounded (given the deviation of the temperature of both policies is bounded), then the divergence between the task-specific policy and the personalised policy can be bounded as stated in Theorem 4.1. This theorem constraints the proposed fusion method to arbitrarily deviate from the task policy.
Theorem 4.1
Let $Q$ and $Q^{\prime}$ be Q-values corresponding to the task policy and the intent-specific policy respectively. Let $\pi_{\psi}(a|s)$ and $\pi_{\phi}(a|s)$ represent the respective policies, with corresponding temperatures $T_{\psi}$ and $T_{\phi}$. Let $\left\lVert Q(s,a)-Q^{\prime}(s,a)\right\rVert_{2}<\epsilon\ \forall s\in\mathcal{S}\ and\ a\in\mathcal{A}$ and $\left\lVert T_{\psi}-T_{\phi}\right\rVert_{2}<\delta$. Then,
$KL\left(\pi_{\phi}(a|s)\middle\|\pi_{f}(a|s)\right)\leq\log\left(Z\right)+\dfrac{1}{2}\left(\dfrac{Q^{*}\delta+\epsilon T_{\phi}}{T_{\phi}T_{\psi}}\right)+\dfrac{1}{2}\log\left(\zeta\right)\ \forall a\in\mathcal{A},s\in\mathcal{S}$,
where $Z$ is the normalising factor of $\pi_{f}$, $Q^{*}=argmax_{a\in\mathcal{A}}Q(s,a)$ and $\zeta=\dfrac{h(Q^{\prime},T_{\psi})}{h(Q,T_{\phi})}$, here $h(Q,T)=\sum_{a}exp\left(\dfrac{Q}{T}\right)$.
Proof
$KL\left(\pi_{\phi}\middle\|\pi_{f}\right)=\sum_{a}\pi_{\phi}\log\left(\dfrac{\pi_{\phi}Z}{\pi_{f}}\right)$
$=\log\left(Z\right)+\sum_{a}\pi_{\phi}\left(\log\left(\sqrt{e^{\dfrac{Q}{T_{\phi}}-\dfrac{Q^{\prime}}{T_{\psi}}}}\right)+\log\left(\sqrt{\zeta}\right)\right)$
$=\log\left(Z\right)+\sum_{a}\dfrac{\pi_{\phi}}{2}\left(\dfrac{Q}{T_{\phi}}-\dfrac{Q^{\prime}}{T_{\psi}}+\log\left(\zeta\right)\right)$
$=\log\left(Z\right)+\sum_{a}\dfrac{\pi_{\phi}}{2}\left(\dfrac{Q}{T_{\phi}}-\dfrac{Q^{\prime}}{T_{\psi}}\right)+\dfrac{1}{2}\log\left(\zeta\right)$
$=\log\left(Z\right)+\sum_{a}\dfrac{\pi_{\phi}}{2}\left(\dfrac{QT_{\psi}-Q^{\prime}T_{\phi}}{T_{\phi}T_{\psi}}\right)+\dfrac{1}{2}\log\left(\zeta\right)$
$=\log\left(Z\right)+\sum_{a}\dfrac{\pi_{\phi}}{2}\left(\dfrac{QT_{\psi}-Q^{\prime}T_{\phi}+QT_{\phi}-QT_{\phi}}{T_{\phi}T_{\psi}}\right)+\dfrac{1}{2}\log\left(\zeta\right)$
$\leq\log\left(Z\right)+\sum_{a}\dfrac{\pi_{\phi}}{2}\left(\dfrac{Q\delta+\epsilon T_{\phi}}{T_{\phi}T_{\psi}}\right)+\dfrac{1}{2}\log\left(\zeta\right)$
$\leq\log\left(Z\right)+\dfrac{1}{2}\left(\dfrac{Q^{*}\delta+\epsilon T_{\phi}}{T_{\phi}T_{\psi}}\right)+\dfrac{1}{2}\log\left(\zeta\right)$
The above theorem establishes the boundedness of the divergence between the task-specific policy and the personalised policy. Nevertheless, we can consider other choices for policy fusion, such as the weighted average or the product of the two policies. However, the former does not operate on the common support, and the latter does not satisfy the invariability constraint, even when the policies are exactly the same, as evidenced in the Lemma 2. For a detailed analysis, we refer the reader to the supplementary Section 0.E.
Lemma 2
Let $\pi_{\psi}$ and $\pi_{\phi}$ be intent-specific policy and task-specific policies respectively and $\pi_{\psi}$ is not a random policy. Then $\pi_{\phi}$ and $\tfrac{1}{Z}\pi_{\psi}\pi_{\phi}$, where $Z$ is normalising factor are not invariant policies in any condition.
Proof
$KL\left(\pi_{\phi}|\pi_{\phi}\pi_{\psi}\right)=$ $\sum_{a}\pi_{\phi}\log\left(\dfrac{\pi_{\phi}Z}{\pi_{\phi}\pi_{\psi}}\right)$
$=$ $\sum_{a}\pi_{\phi}\log\left(\dfrac{Z}{\pi_{\psi}}\right)$
$=$ $\log\left(Z\right)-\sum_{a}\pi_{\phi}\log\pi_{\psi}$
Where $Z$ is the normalising factor, $Z=\sum_{a}\pi_{\phi}\pi_{\psi}$. Assume the expression above is 00 $\Rightarrow\log\left(Z\right)=\sum_{a}\pi_{\phi}\log\pi_{\psi}$
$\log\left(\sum_{a}\pi_{\phi}\pi_{\psi}\right)=\sum_{a}\pi_{\phi}\log\pi_{\psi}$
$\Rightarrow\log\left(E_{\pi_{\phi}}\left(\pi_{\psi}\right)\right)=E_{\pi_{\phi}}\left(\log\left(\pi_{\psi}\right)\right)$
This expression is valid only if $\pi_{\phi}$ is a random policy, which is a contradiction.
Despite the compliance to invariance constraints of the fused policy, in practice, we found that this approach of statically fusing policies is subject to certain pitfalls, as explained in the next section. Subsequently, we address the issue through our novel dynamic policy fusion approach in Section 4.5.
4.4 Pitfalls of Static Fusion
With the static policy fusion technique described in Section 4.3, a potential challenge arises wherein one of the component policies over-dominates the other. Consequently, the agent may disproportionately exhibit the corresponding behaviour, resulting in noncompliance with either the task-specific policy or the intent-specific policy.
To illustrate this phenomenon, we consider a 2D Navigation scenario where an agent is tasked with reaching a target location. However, a human may wish for some checkpoint state, different from the target state, to be visited before the agent reaches its target. In this case, the task-specific policy corresponds to the actions along the shortest path towards the target, and an intent-specific policy would correspond to one inferred from human feedback, which exclusively favours visiting the checkpoint state. A static policy fusion approach in this case could lead to over-dominance of one of the component policies. For example, over-dominance of the intent-specific policy would lead to the agent visiting the checkpoint state indefinitely. This motivates the need for fusing the policies in a dynamic fashion, such that the agent respects the human’s preferences, while also simultaneously completing the task at hand. We develop such a dynamic fusion technique to control the relative dominance of the individual policies by modulating the temperature parameter $T_{\psi}$ of the intent-specific policy. We now describe the details of our dynamic fusion strategy.
Figure 2: Plot showing the variation of $T_{\psi}$. The temperature rises when the accumulated $r^{\prime}$ reaches $\eta$.
4.5 Dynamically Modulating Policy Dominance
To mitigate unintended policy dominance, we adopt the idea that when the temperature parameter $T_{\psi}$ in Equation (3) is increased, the probability distribution of actions tends to become uniform, thereby reducing the influence of the intent-specific policy on the personalised policy. We therefore modulate $T_{\psi}$ depending on whether the fused personalised policy exhibits over-adherence or under-adherence to the intent-specific policy $\pi_{\psi}$.
When to increase $T_{\psi}$? $T_{\psi}$ should increase if the intent-specific policy $\pi_{\psi}$ is enforced too strongly, which is characterised by high accumulated human-induced rewards. We design $T_{\psi}$ to increase when the accumulated human-induced reward surpasses an accumulated reward threshold $\eta$.
However, we note that human intent can be specified in various modes: preference (where a state is preferred over others), avoidance (where the preference is to avoid a particular state) or mixed (where the preference is to avoid certain states and to prefer certain others). In avoidance cases, human feedback assigns lower scores to unfavourable trajectories (those that pass through the state(s) to be avoided) while in the preference case, higher scores are assigned to preferred trajectories (those that pass through the preferred state(s)). Depending on the trajectory score from the human, the LSTM assigns rewards to each state-action pair as in Equation (1) which we refer to as human-induced rewards.
To modulate $T_{\psi}$, we first compute the shifted rewards as follows:
$r^{\prime}=r-mean(r),$ (5)
where $r$ is the vector that contains the rewards for different actions from a given state. The purpose of this shifting operation is to ensure that irrespective of the mode of human intent, the reward vector $r^{\prime}$ contains elements with both positive as well as negative values. This shifted human-induced reward $r^{\prime}$ is then used to modulate $T_{\psi}$ as:
$T_{\psi}(t)=\max\left(T_{min},\frac{T_{max}}{1+\exp\left(-\left(g(t)-\eta\right)\right)}\right),$ (6)
where $T_{min}$ and $T_{max}$ are the minimum and maximum allowable temperatures and $g(t)=\sum_{t^{\prime}=0}^{t}r^{\prime}(s_{t^{\prime}},a_{t^{\prime}})$ is the accumulated shifted human-induced reward. A plot of $T_{\psi}$ is shown in the Figure 2.
How does shifting (Equation (5)) enable two-way switching? A high accumulated shifted human-induced reward ($>\eta$) indicates the agent has obeyed the intent-specific policy $\pi_{\psi}$ disproportionately more, which flags the need to mitigate the influence of $\pi_{\psi}$. Since the temperature in Equation (6) is monotonic with respect to accumulated reward $r^{\prime}$, a higher accumulated $r^{\prime}$ outputs a higher temperature, and this reduces the influence of the $\pi_{\psi}$ in the policy fusion step. Conversely, if the accumulated reward $r^{\prime}$ $<\eta$, the temperature is lowered, which strengthens $\pi_{\psi}$. Therefore, this setup enables a two-way dynamic switching between the policies. Based on the threshold $\eta$, the influence of $\pi_{\psi}$ is reduced when the accumulated $r^{\prime}$ is high, and strengthened when it is too low. Our overall algorithm for personalisation is summarised in Algorithm 1.
Algorithm 1 Personalising the action selection in an episode
Input: DQN and LSTM Networks $\phi$ and $\psi$, accumulated reward threshold $\eta$, Task-specific policy temperature $T_{\phi}$, Min and Max temperature $T_{min}$ and $T_{max}$, slope of the sigmoid $m$
Reset the environment to get the initial state $s_{0}$
Initialise: $Q^{\prime}(s_{0},a_{0})=0$, $g(0)=0$, $T_{\psi}=max(T_{min},\frac{T_{max}}{1+exp(m\times\eta))})$, $done=False$, $t=0$
1:while not $done$ do
2:     $Q_{t}\leftarrow[]$triangleright DQN Q-values reset the vector to null
3:     $Q^{\prime}_{t}\leftarrow[]$triangleright LSTM Q-values reset the vector to null
4:     $r_{t}\leftarrow[]$triangleright Human-induced reward reset the vector to null
5:     $Q_{t}\leftarrow\phi(s_{t})$triangleright Invoking DQN
6:     $Q^{\prime}_{t}\leftarrow\phi(s_{t},a)$triangleright Invoking LSTM for each action
7:     Update $r_{t}$ as in Equation (1)
8:     $\pi_{\phi}\leftarrow BoltzmanDistribution(Q_{t},T_{\phi})$triangleright Task-specific policy
9:     $\pi_{\psi}\leftarrow BoltzmanDistribution(Q^{\prime}_{t},T_{\psi})$triangleright Intent-specific policy
10:     $a_{t}\leftarrow\arg\max_{a\in\mathcal{A}}\sqrt{\pi_{\phi}(a|s_{t})\times\pi_{\psi}(a|s_{t})}$triangleright Policy fusion
11:     $r^{\prime}\leftarrow r_{t}-mean(r_{t})$triangleright Adjusting as in Eq (5)
12:     $g(t)\leftarrow\sum_{t^{\prime}=0}^{t}r^{\prime}(s_{t^{\prime}},a_{t^{\prime}})$triangleright Accumulated shifted human-induced reward
13:     $T\leftarrow max(T_{min},\frac{T_{max}}{1+exp(-m(g-\eta))})$triangleright Updating $T$
14:     Execute the action $a_{t}$ from state $s_{t}$ to transition to $s_{t+1}$
15:     $t\leftarrow t+1$triangleright Increment the time-step
16:     if Episode terminate then
17:         $done=True$.
18:     end if
19:end while
5 Experiments
We demonstrate the effectiveness of our approach in our custom 2D Navigation environment, Highway [highwayenv], Pong [26] depicted in Figure 3. Specifically, we conduct personalisation experiments in three scenarios: Avoidance, Preference and Mixed  [^1] Code for our implementation can be found here. For Pong, we only consider the Preference scenario. To learn the intent-specific policy, we sampled 2000 trajectories from saved trajectory data. Now we describe the environments as follows:
Figure 3: Snippets from the 2D environment (a), Pong (b) and Highway (c). In 2D navigation, the observation is converted to a grey-scale image. In Pong (b), the right paddle is controlled by the CPU and the left is controlled by the agent. The preferred region corresponds to the centre, as indicated. In the Highway environment (c), lanes marked with green and red boxes represent desired and undesired lanes respectively.
2D Navigation: This is our custom environment where an agent is tasked with navigating to a target state while avoiding undesired states and visiting desired ones. The observation consists of a $40\times 40$ grayscale image with four directional movements allowed. Actions that put the agent out of the frame are invalid. A +1 reward is granted upon reaching the target location, and a 0 reward is given for all other actions. The episode concludes within 20 timesteps or upon reaching the target.
Pong: In Pong, the observations correspond to ball x (horizontal) and y (vertical) position and velocity, player y position and velocity and CPU y position. Actions involve increasing, decreasing, or changing the velocity of the paddle. We have changed the original environment reward scheme and the termination criteria as follows. A +1 reward is obtained upon winning, a 0 reward for losing the game and 0.5 if the trajectory terminates. Pong concludes after 1000 timesteps or upon game outcome (win or loss).
Highway: The goal is to navigate through traffic as fast as possible. The observations consist of a 26-dimensional vector, horizontal (x) and vertical(y) coordinates, corresponding x and y-velocities of five nearby vehicles, and the current lane. The agent can change lanes, stay idle, or adjust speed (move faster or slower). Any action that takes the agent out of frame is void. This environment has a dense reward setting, with positive rewards granted for maximizing speed, and negative rewards incurred upon collision with other vehicles. These rewards are normalized between 0 and 1. Episode termination occurs after 50 timesteps or upon collision with other vehicles.
Baselines: We compare our proposed method against 4 baselines.
  • DQN [16]: which learns a policy that focuses solely on the task objective.
  • RUDDER [1]: which learns a policy focusing on just the human objective.
  • Static: Static fusion method described in Section 4.3 designed to achieve both objectives.
  • MORL [18]: A multi-objective RL baseline scalarises the vector reward of task and human-induced reward by linearisation.
We refer the reader to the supplementary material for details of the hyperparameters used and baseline implementations.
Table 1 summarises the different environments with corresponding human and task objectives. To personalise the behaviour, we assume human users have lane preferences or aversions in Highway. Likewise, in 2D Navigation, humans prefer the agent to avoid or visit certain regions. In Pong, personalisation is introduced by favouring specific paddle positions.
Table 1: Table showing different personalisation modes and corresponding human and task objectives.
Env Mode Human objective Task objective 2D Navigation Preference Prefer a region Move to the target Avoidance Avoid a region Mixed Prefer a region and avoid another Pong Preference Prefer certain paddle positions Win the game Highway Preference Prefer a lane Move as fast as possible Avoidance Avoid a lane without colliding Mixed Prefer a lane and avoid another with other vehicles
Although Section 4 discussed actions being sampled from policies, in practical applications, it is common to exploit greedy actions once the policy is learned. Hence, in all our experiments, we choose actions as $argmax_{a\in\mathcal{A}}\sqrt{\pi_{\phi}(a|s_{t})\times\pi_{\psi}(a|s_{t})}$. This is akin to using a low-temperature parameter setting in the fused Boltzmann policy.
Table 2: Results in 2D Navigation. Undesired/Desired region is the average visit out of a total of 20 time steps. Values marked in blue indicate a top 15% task performance score relative to the best baseline, and those in red indicate a top 15% measure relative to the baseline that best adheres to human preferences. The up arrow $\uparrow$ indicates higher is better, and the down arrow $\downarrow$ indicates lower is better. The result averaged over 10 seeds with each seed containing 300 episodes with $T_{\phi}=0.4$, $T_{min}=1$, $T_{max}=10$, and $\eta=0$.
Mode Method Desired region $\uparrow$ Undesired region $\downarrow$ Score $\uparrow$ Preference DQN $0.085\pm 0.005$ $1.000\pm 0.000$ RUDDER $6.546\pm 0.342$ $0.006\pm 0.005$ MORL $9.575\pm 0.021$ $000.000\pm 0$ 0.000 Static $3.788\pm 0.452$ $0.602\pm 0.048$ Dynamic (Ours) $1.459\pm 0.140$ $1.000\pm 0.000$ Mixed DQN $0.050\pm 0.003$ $0.183\pm 0.006$ $1.000\pm 0.000$ RUDDER $2.790\pm 0.311$ $000.000\pm 0$ 0.000 $0.007\pm 0.007$ MORL $0.192\pm 0.022$ $0.119\pm 0.016$ $0.396\pm 0.035$ Static $1.116\pm 0.318$ $000.000\pm 0$ 0.000 $0.894\pm 0.035$ Dynamic (Ours) $0.268\pm 0.058$ $000.000\pm 0$ 0.000 $1.000\pm 0.000$ Avoidance DQN - $0.094\pm 0.006$ $1.000\pm 0.000$ RUDDER - $000.000\pm 0$ 0.000 $0.292\pm 0.055$ MORL - $0.026\pm 0.021$ $0.894\pm 0.099$ Static - $0.005\pm 0.005$ $1.000\pm 0.000$ Dynamic (Ours) - $0.006\pm 0.005$ $1.000\pm 0.000$
5.1 Results
We first demonstrate the pitfalls of static fusion as described in Section 4.4 and compare it against dynamic fusion in 2D Navigation. As the performance of the static fusion baseline is sensitive to $T_{\psi}$, we tested a range of temperature values and found $T_{\psi}=T_{max}/2$ as the best-performing agent across all three modes. As observed in Table 2, such a static fusion agent performed well in the Avoidance mode, exhibiting behaviours that avoid the undesired region, while achieving high scores. However, as seen in the Preference and Mixed modes, the static fusion agent incurs severe drops in its score and instead exhibits a high number of visits to the desired region. This is indicative of the over-dominance of the intent-specific policy component in the personalised policy. We confirmed this by examining the agent’s behaviour, which was found to repeatedly visit the desired region, ignoring the primary navigation task. Although the performance of static fusion could possibly be improved in individual modes of personalisation, it would require extensive tuning of the temperature hyperparameter, which renders this approach practically infeasible.
Regarding the other baselines, as expected, DQN prioritised task performance (Score), while RUDDER exhibited over-adherence to human preferences at the cost of task performance. The MORL baseline which is designed to balance multiple objectives was found to be less effective compared to our proposed dynamic fusion method in terms of both task and human objectives. We attribute this under-performance to an inherent issue in MORL, where it struggles to determine which objective to prioritise, particularly in situations where conflicts in action selection arise. This issue becomes apparent in the mixed case of 2D Navigation, where MORL attempts to find a global solution satisfying both objectives, whereas our dynamic policy fusion method enables policy corrections locally, wherever adherence to human preferences is deemed necessary. Overall, the dynamic fusion agent strikes a balance between the primary task and intent-specific policy, assimilating both aspects into the behaviour, irrespective of the mode of personalisation.
These results establish fundamental limitations of static fusion and other baselines and highlight the need for dynamic fusion in order to satisfy both human and task objectives.
We further examine the properties of dynamic fusion in the Highway and Pong environments in Tables 3 and 4 respectively with the results averaged over 10 seeds each with 20 episodes. Here, DQN and RUDDER perform as expected, prioritising the task and human objectives respectively. In Highway, as seen in Table 3, DQN maximises the score and minimises hits (collisions), whereas RUDDER exhibits behaviours avoiding the undesired (first) lane and favouring the desired (third) lane. Similarly, as indicated in Table 4 in Pong, while DQN maximises the task score at the cost of seldom visiting the desired region, RUDDER causes the agent to remain in the preferred region, consistent with human preference (but incurring low scores). When compared to the dynamic policy fusion, the performance of the MORL baseline falls short of achieving both objectives. This is evident from the higher frequency of undesired lane visits in both the Avoidance and Mixed modes, and the lower frequency of desired lane visits in the Preference and Mixed modes. These results are detailed in Table 3. A similar trend is observed in Pong, where the MORL baseline not only failed to score but also spent less time in the preferred region compared to our method. These results indicate that dynamic policy fusion, while sometimes incurring a small reduction in the optimal score (achieved by DQN), enables the agent to adhere to the required preference (be it in the Avoiding, Preference or Mixed modes) while completing the primary task at hand. Additional experiments on different $\alpha$ in the Highway environment are provided in supplementary Section 0.G.
Table 3: Results for the Highway environment. Undesired/Desired lane columns indicate the average number of visits in 50 time steps. Values marked in blue indicate a top 15% task performance score relative to the best baseline, and those in red indicate a top 15% measure relative to the baseline that best adheres to human preferences. $\uparrow$ indicates higher values are better and $\downarrow$ indicates lower values are better for the quantity specified in the column. The hyper-parameters chosen are $T_{\phi}=0.6,T_{max}=5,T_{min}=0.3,\eta=0$.
Mode Method Desired lane $\uparrow$ Undesired lane $\downarrow$ Hits $\downarrow$ Score $\uparrow$ Avoidance DQN - $10.59\pm 0.96$ $0.11\pm 0.01$ $39.59\pm 0.46$ RUDDER - $00.00\pm 0$ 0.00 $0.73\pm 0.11$ $18.43\pm 2.54$ MORL - $5.80\pm 1.07$ $0.14\pm 0.23$ $37.44\pm 0.52$ Dynamic (Ours) - $0.19\pm 0.11$ $0.09\pm 0.02$ $38.80\pm 0.51$ Preference DQN $10.9\pm 0.49$ $0.07\pm 0.02$ $40.25\pm 0.64$ RUDDER $31.84\pm 3.45$ $0.30\pm 0.06$ $28.45\pm 1.65$ MORL $11.58\pm 1.12$ $0.07\pm 0.01$ $38.14\pm 0.30$ Dynamic (Ours) $29.91\pm 1.38$ $0.06\pm 0.01$ $39.27\pm 0.59$ Mixed DQN $11.83\pm 0.64$ $9.59\pm 0.82$ $0.08\pm 0.02$ $40.06\pm 0.79$ RUDDER $26.26\pm 5.90$ $00.00\pm 0$ 0.00 $0.50\pm 0.12$ $24.67\pm 3.00$ MORL $12.61\pm 1.20$ $12.36\pm 1.58$ $0.07\pm 0.01$ $36.97\pm 0.59$ Dynamic (Ours) $25.21\pm 2.55$ $0.27\pm 0.06$ $0.07\pm 0.04$ $39.15\pm 0.49$
Table 4: Table showing the result in the Pong environment with $\eta=0$. Values marked in blue indicate a top 15% task performance score relative to the best baseline, and those in red indicate a top 15% measure relative to the baseline that best adheres to human preferences. % Desired region is the average fraction (in percentage) of an episode the agent stayed within the desired region shown in Figure 3 (b).
Method % Desired region $\uparrow$ Score $\uparrow$ DQN $8.00\pm 0.63$ $0.57\pm 0.01$ RUDDER $89.99\pm 0.14$ $00.00\pm 0$ 0.00 MORL $41.68\pm 0.02$ $00.00\pm 0$ 0.00 Dynamic (Ours) $55.60\pm 4.48$ $0.48\pm 0.02$
Table 5: Results correspond to the Highway environment with different values of $\eta$. $\uparrow$ indicates higher values are better and $\downarrow$ indicates lower values are better for the quantity specified in the column. The hyper-parameters chosen are $T_{\phi}=0.6,T_{max}=5,T_{min}=0.3$.
Mode $\eta$ Desired lane $\uparrow$ Undesired lane $\downarrow$ Hits $\downarrow$ Score $\uparrow$ Avoidance 0 - $0.28\pm 0.08$ $0.06\pm 0.01$ $39.83\pm 0.59$ 1 - $0.35\pm 0.12$ $0.08\pm 0.02$ $38.96\pm 0.87$ 2 - $0.19\pm 0.11$ $0.09\pm 0.02$ $38.80\pm 0.51$ Preference 0 $25.14\pm 1.44$ $0.05\pm 0.02$ $40.02\pm 0.61$ 1 $24.90\pm 1.52$ $0.10\pm 0.03$ $38.84\pm 0.84$ 2 $29.91\pm 1.38$ $0.06\pm 0.01$ $39.27\pm 0.59$ Mixed 0 $21.52\pm 2.00$ $0.85\pm 0.33$ $0.12\pm 0.01$ $38.50\pm 0.43$ 1 $23.96\pm 1.67$ $0.56\pm 0.13$ $0.07\pm 0.01$ $39.56\pm 0.45$ 2 $25.21\pm 2.55$ $0.27\pm 0.06$ $0.07\pm 0.04$ $39.15\pm 0.49$
Table 6: Table depicts the sensitivity analysis of $T_{max}$ in Highway environment mixed case. The hyper-parameters chosen are $T_{\phi}=0.6,T_{min}=0.3$ and $\eta=0$. $\uparrow$ indicates higher values are better and $\downarrow$ indicates lower values are better for the quantity specified in the column.
[…]
Effect of $\eta$: In Table 5, we examine the effect of $\eta$ on the performance of an agent under dynamic fusion. In theory, a higher $\eta$ would make it more likely for $T_{\psi}$ to remain low (refer to Figure 2), and thus the intent-specific policy becomes sharper. So while fusing the policies together, the obtained personalised policy become dominated by intent-specific policy. Our observations in Table 3 are in agreement with this behaviour, as in all preference modes, as $\eta$ increases, the desired lane visit increases and the undesired lane visit decreases, suggesting that personalised policy is dominated by the intent-specific policy component. In all of our experiments, we choose $\eta=0$. However, an ideal choice of $\eta$ may deviate from this value in other environments.
Effect of $T_{max}$: The parameter $T_{max}$ represents the maximum temperature in Equation (6). Theoretically, as $T_{max}$ is increased the intent-specific policy should weaken. The observation in Table 6 supports this notion that when $T_{max}$ is increased, the desired lane visit decreases and the undesired lane visit increases suggesting a declining influence of intent-specific policy.
6 Limitations and Future Work
Even though our proposed dynamic policy fusion approach dynamically optimises two objectives, in its current state, it is restricted to discrete action spaces. Extension to continuous action domains could be explored in future work. Additionally, the choice of hyper-parameters is done using trial and error; reducing or learning the number of parameters will enhance the utility of the method. Our method does not capture the varying preferences of the human. To address this limitation, future research could explore adaptive methods, perhaps based on meta-RL [3] to adjust to changing human preferences with a few shot learning of the LSTM network. The experiments presented in this work also make use of a noise-free ground truth function to simulate human feedback. We acknowledge that in reality, human feedback may be subject to noise, which could affect the robustness of our presented approach. Further research is necessary to adapt our approach to be more robust to noise, the performance of which could then be evaluated with noisy feedback from real humans. Another practical aspect to consider is that the number of trajectories one could expect a human to label is limited. From an initial investigation, our approach does exhibit a drop in performance as fewer labelled trajectories become available. Although this problem could partially be mitigated through the use of semi-supervised methods [29] to automatically generate trajectory score labels, we recognise that more noise-tolerant and feedback-efficient variants of our method need to be explored. Nevertheless, we believe the proposed dynamic policy fusion method presents a novel and an in-principle, practical approach for adapting and personalising already trained policies.
7 Conclusion
We proposed a novel approach to personalise a trained policy to respect human-specified preferences using trajectory-level human feedback, without any additional interactions with the environment. Using an LSTM-based approach to infer human intent, we designed a theoretically grounded dynamic policy fusion method to ensure the resulting policy completes a given task while also respecting human preferences. We empirically evaluated our approach on the Highway, 2D Navigation and Pong environments and demonstrated that our approach is capable of handling various modes of intent while only minimally compromising the task performance. We believe our approach presents an elegant and scalable solution to the problem of personalising pretrained policies. With a growing focus on personalisation in applications such as chatbots, robotic assistants, self-driving vehicles, etc., we believe our approach has the potential for imminent and widespread impact.
Appendix 0.A Simulating human feedback
Our method involves collecting human feedback for trajectories used for training the task-specific policy. This is done by simulating humans. We conducted experiments in three personalisation modes such as preference, Avoidance and Mixed. In each case, we counted the number of times the agent met the personalisation criteria within a trajectory and this is regarded as the ground truth score for a trajectory. For instance, In 2D navigation, if the human wishes the agent to visit a preferred region, we count the number of times the agent visited that state and a positive score is given to that trajectory. Similarly, in the avoidance case, a negative score is given and in the mixed case, we sum both the positives and the negatives.
Applying this method in a real-world context could involve several challenges. Using a real human for scoring trajectories, for example, may have limitations with regards to the number of trajectories to be labeled. Fewer labeled trajectories may have an adverse effect on the performance of our approach. In addition to this, a real subject would be prone to fatigue and may provide more arbitrary score labels. For instance, their scores may be biased based on the quality of trajectories presented to them immediately prior to scoring. These, along with a number of other challenges may need to be overcome before applying this approach to real-world scenarios.
Appendix 0.B Training of LSTM
As described in the main paper, we used the same trajectory training data of the task-specific policy to train the LSTM. We sampled 2000 trajectories along with the human feedback as the training data.
The input to the LSTM is the state and the action chosen at time $t$ and it outputs the corresponding Q-values. We made a custom one-hot encoding of the action to nearly match the dimension of the state vector. For example, in the Highway environment, the state vector is 26 dimensional and there are 5 actions. We encode the selected action to a 25-dimensional vector with one filling in the corresponding 5 positions of the action and the rest being zero. We then condition the state vector with the action vector and train the LSTM. An overview of this dataflow is outlined in Figure 4.
Figure 4: LSTM network training overview. The state vector and the action vector are fed to the FiLM network. The FiLM network conditions the action vector with a state vector and this conditioned vector is the input to the LSTM. LSTM outputs the estimated q-value.
The LSTM network architecture is single-layered with 64 units. The output gate and forgot gate of the LSTM are set to 1 as in [1] as we don’t want to forget anything while performing the credit assignment. The optimiser details are provided in Table 7. We used the same loss functions as that of [1] to train the LSTM.
$L_{m}=(l-\tilde{Q}_{H})^{2},$ (7)
$L_{c}=\frac{1}{H+1}\sum_{t=0}^{H}(l-\tilde{Q}_{t})^{2},$ (8)
$L_{e}=\frac{1}{H-\delta+1}\sum_{t=0}^{H-\delta}(\tilde{Q}_{t+\delta}-\beta)^{2},$ (9)
where $l$ is the human label and $\tilde{Q}_{t}$ is the LSTM output at timestep $t$ and $H$ is the final timestep. $\beta$ is the prediction of $\tilde{Q}_{t+\delta}$ at timestep $t$. $\delta$ is the forward lookup which is set to 3 in all our experiments.
The final loss function is,
$L=L_{m}+\frac{1}{10}(L_{c}+L_{e}).$ (10)
The $L_{c}$ term aims to allocate full credit to each time step without considering its impact on future steps. Basically, it constrains the expected future reward to zero. To get $\beta$ in $L_{e}$, we added a separate head to the LSTM network which predicts the Q-values at $t+\delta$ at time $t$. Both $L_{e}$ and $L_{c}$ act as auxiliary loss functions for reward redistribution.
Table 7: Table showing the LSTM Optimizer details
Particulars Value Learning rate $10^{-4}$ Weight decay $10^{-8}$ Gradient clipping $10$ Optimizer ADAM
Table 8: Table showing the VAE Optimizer details
Particulars Value Learning rate $10^{-3}$ Weight decay $10^{-6}$ Gradient clipping $10$ Optimizer ADAM
Table 9: VAE encoder architecture
Conv layer Input depth Output Depth Kernel size Stride Padding Conv 1 1 16 3 2 0 Conv 2 16 32 3 2 0 Conv 3 32 64 3 1 0 Conv 4 64 128 3 1 0 Conv 5 128 256 3 1 1
Table 10: VAE decoder architecture
ConvT layer Input depth Output Depth Kernel size Stride Padding ConvT 1 256 128 3 1 1 ConvT 2 128 64 3 1 0 ConvT 3 64 32 3 2 0 ConvT 4 32 16 5 1 0 ConvT 5 16 1 4 2 0
Experiments were conducted across three distinct environments, each employing different modes of personalisation. For the 2D Navigation and Highway environments, we implemented experiments in avoidance, preference, and mixed modes. However, for the Pong environment, personalisation was conducted in preference mode. The LSTM loss function for all modes across the 2D Navigation, Highway, and Pong environments is depicted in Figures 5, 6, and 7, respectively.
Given that the horizon for the Pong environment is 1000, which is relatively long, we opted to train the LSTM using snippets of the trajectory that contain 50 timesteps. We simulated the feedback for these short snippets. For all other environments, we utilized the full-length trajectory in our training process.
(a) Avoidance
(b) Preference
(c) Mixed
Figure 5: LSTM loss function of 2D navigation across all modes of personalisation
(a) Avoidance
(b) Preference
(c) Mixed
Figure 6: LSTM loss function of Highway across all modes of personalisation
Figure 7: LSTM loss function of Pong
Figure 8: VAE loss function. The scale of the loss is high because the loss is summed over the data for each mini batch
Appendix 0.C Training of VAE
VAE is pretrained and only used if the observation of the environment is pixel-based. Training data is the same as the training data of the task policy. We use the conventional VAE loss functions which consist of reconstruction loss and KL-divergence loss. The encoder and decoder architectures are tabulated in Tables 9, 10 respectively. A fully connected linear layer is used to map the output from the encoder to a dimension of 512 subsequently the output is mapped to a latent dimension (set as 16) using another linear layer. Optimiser details of the VAE is shown in Table 8. The loss curve is shown in Figure 8.
Appendix 0.D Different choices of static fusion
Policy fusion is a process of combining two or more policies to produce a new policy. [24] studied various policy fusion methods which are as follows.
0.D.1 Product Policy
The fused policy is obtained by multiplying different policies. The action selection process from the $\pi_{f}$ is as follows:
$\arg\max_{a\in\mathcal{A}}\pi_{\phi}(a|s_{t})\times\pi_{\psi}(a|s_{t})$ (11)
0.D.2 Mixture Policy
In this fusion method, the fused policy is obtained by taking the average of the policies. The action selection process is expressed as follows:
$\arg\max_{a\in\mathcal{A}}\frac{\pi_{\phi}(a|s_{t})+\pi_{\psi}(a|s_{t})}{2}$ (12)
0.D.3 Entropy Threshold
This method selects the action based on the entropy of each policy at state $s_{t}$. Let $\mathcal{H}{\phi}(s_{t})$ represent the entropy of the task policy, and $\mathcal{H}{\psi}(s_{t})$ represent the entropy of the Intent-specific policy at state $s_{t}$. The action is selected according to the following conditions:
$\begin{cases}\arg\max_{a\in\mathcal{A}}\pi_{\psi}(a|s_{t})&\text{If }\mathcal{H}{\psi}(s_{t})<\mathcal{H}{\phi}(s_{t})+\epsilon\\ \arg\max_{a\in\mathcal{A}}\pi_{\phi}(a|s_{t})&Otherwise\end{cases}$ (13)
Here, $\epsilon$ is a small positive value.
0.D.4 Entropy Weighted
This fusion method computes a weighted average of the two policies, where the weights are determined by the minimum entropy among the two policies. Let $\mathcal{H}^{*}=\min\left(\mathcal{H}{\psi}(s_{t}),\mathcal{H}{\phi}(s_{t})\right)$. The action is selected as follows:
$\arg\max_{a\in\mathcal{A}}\left(\mathcal{H}^{*}\times\pi_{\phi}(a|s_{t})+(1-\mathcal{H}^{*})\times\pi_{\psi}(a|s_{t})\right)$ (14)
The term ×$\mathcal{H}^{*}\times\pi_{\phi}(a|s_{t})$ represents the weighted contribution of the task policy, and $(1-\mathcal{H}^{*})\times\pi_{\psi}(a|s_{t})$ represents the weighted contribution of the Intent-specific policy.
Appendix 0.E Analysis of other choices of policy fusion
In this section, we focus on analyzing two common fusion methods: the weighted average and product fusion methods through the lens of personalisation. The fusion strategy should comply with two constraints as defined in the main paper. Firstly, the fused policy should act on the common support of the policies being fused and secondly, satisfy the invariability constraint.
The former method is defined as $\tfrac{1}{Z}\alpha\pi_{1}+(1-\alpha)\pi_{2}$, where $\pi_{1}$ and $\pi_{2}$ are the policies to be fused, $Z$ is normalising factor and $\alpha$ is a weight parameter. This method satisfies the invariability property, ensuring that the fused policy remains unchanged when both input policies are similar. However, this fusion strategy acts on the union of the support of the input policies, meaning that actions are selected to maximize either one objective and not both simultaneously.
The latter method involves taking the product of the two policies: $\tfrac{1}{Z}\pi_{f}(a|s)=\pi_{1}(a|s)\times\pi_{2}(a|s)$. While this fusion operates on the intersection of the support of the input policies, ensuring that actions are selected to maximize both objectives simultaneously, it fails to satisfy the invariability property as shown in Lemma 2,
However, the fused policies produced by the product can be bounded as captured by the following theorem,
Theorem 0.E.1
Let $Q$ and $Q^{\prime}$ be Q-values corresponding to the task policy and the intent specific policy respectively. Let $\pi_{\psi}(a|s)$ and $\pi_{\phi}(a|s)$ represent the respective policies, with corresponding temperatures $T_{\psi}$ and $T_{\phi}$. Let $\left\lVert Q(s,a)-Q^{\prime}(s,a)\right\rVert_{2}<\epsilon\ \forall s\in\mathcal{S}\ and\ a\in\mathcal{A}$ and $\left\lVert T_{\psi}-T_{\phi}\right\rVert_{2}<\delta$. Then,
$KL\left(\pi_{\phi}(a|s)\middle\|\pi_{f}(a|s)\right)\leq\log\left(Z\right)+\left(\dfrac{Q^{*}\delta+\epsilon T_{\phi}}{T_{\phi}T_{\psi}}\right)+\log\left(\zeta\right)\ \forall a\in\mathcal{A},s\in\mathcal{S}$,
where $Z$ is its normalising factor of $\pi_{f}(a|s)$, $Q^{*}=argmax_{a\in\mathcal{A}}Q(s,a)$ and $\zeta=\dfrac{h(Q^{\prime},T_{\psi})}{h(Q,T_{\phi})}$, here $h(Q,T)=\sum_{a}exp\left(\dfrac{Q}{T}\right)$.
Proof
$KL\left(\pi_{\phi}\middle\|\pi_{f}\right)=\sum_{a}\pi_{\phi}\log\left(\dfrac{\pi_{\phi}Z}{\pi_{f}}\right)$
$=\log\left(Z\right)+\sum_{a}\pi_{\phi}\left(\log\left(e^{\dfrac{Q}{T_{\phi}}-\dfrac{Q^{\prime}}{T_{\psi}}}\right)+\log\left(\zeta\right)\right)$
$=\log\left(Z\right)+\sum_{a}\pi_{\phi}\left(\dfrac{Q}{T_{\phi}}-\dfrac{Q^{\prime}}{T_{\psi}}+\log\left(\zeta\right)\right)$
$=\log\left(Z\right)+\sum_{a}\pi_{\phi}\left(\dfrac{Q}{T_{\phi}}-\dfrac{Q^{\prime}}{T_{\psi}}\right)+\log\left(\zeta\right)$
$=\log\left(Z\right)+\sum_{a}\pi_{\phi}\left(\dfrac{QT_{\psi}-Q^{\prime}T_{\phi}}{T_{\phi}T_{\psi}}\right)+\log\left(\zeta\right)$
$=\log\left(Z\right)+\sum_{a}\pi_{\phi}\left(\dfrac{QT_{\psi}-Q^{\prime}T_{\phi}+QT_{\phi}-QT_{\phi}}{T_{\phi}T_{\psi}}\right)+\log\left(\zeta\right)$
$\leq\log\left(Z\right)+\sum_{a}\pi_{\phi}\left(\dfrac{Q\delta+\epsilon T_{\phi}}{T_{\phi}T_{\psi}}\right)+\log\left(\zeta\right)$
$\leq\log\left(Z\right)+\left(\dfrac{Q^{*}\delta+\epsilon T_{\phi}}{T_{\phi}T_{\psi}}\right)+\log\left(\zeta\right)$
Appendix 0.F MORL baseline Implementation Details
To implement the MORL baseline, we construct the reward as a convex sum of the rewards of both objectives as $r=\alpha r_{e}+(1-\alpha)r_{h}$, where $r_{e}$ is the reward obtained from the environment and the $r_{h}$ is the reward obtained from the LSTM as in Equation (1). To match with our zero-shot approach, we change the reward of the transitions of the trajectories saved for the training of LSTM. We then populated the new transitions in the replay buffer and trained a new DQN. We choose $\alpha=0.5$ in all our experiments to balance both objective and normalised the human reward components between $[-1,1]$.
Appendix 0.G Additional experiments
In this section, we summarise some additional experiments conducted. Firstly, we checked the behaviour of the static fusion when the LSTM temperature ($T_{\psi}$) is kept as $T_{min}$ in 2D Navigation environment. In this experiment, we found that the fused policy is dominated by human objectives as shown in Table 12. This is because a low temperature results in a sharper Boltzmann probability distribution. Therefore the intent-policy dominates over the other.
Secondly, We investigated the behaviour change with different reward component weighting ($\alpha$) in the MORL baseline. Since MORL, tries to optimise different objectives simultaneously, the weighting of the reward component plays a significant role. If the weighting of the reward component is high, the objective corresponding to that reward is highly prioritised. We empirically tested this in the Highway environment with weightings as $\alpha=0.3,0.5$ and $0.7$. As the value of the $\alpha$ increases, the policy obtained from MORL shifts more towards the human objective as in Table 13.
Appendix 0.H DQN Training
DQN was utilised to obtain the task-specific policy. For the 2D navigation environment, which involves image-based observations, we adopted the CNN network as detailed in Table 11. In the case of the Pong and Highway environments, an MLP was implemented, featuring three hidden layers with sizes of 256, 512, and 512. We use the ReLU activation function after each layer. We set the minimum epsilon at 0.10, with a decay rate of 0.995. The reward curve for each environment is depicted in Figure 9.
(a) 2D Navigation
(b) Highway
(c) Pong
Figure 9: Smoothed reward curve of DQN for 2D Navigation, Highway and Pong
Table 11: Table showing the CNN architecture of DQN
Particulars Value Channel 32,64,64 Kernel size 11,5,3 Stride 2,2,1 Padding 0,0,0 FC hidden layer 512
Table 12: Result of static fusion when $T_{\psi}=T_{min}$ in 2D Navigation. The up arrow $\uparrow$ indicates higher is better, and the down arrow $\downarrow$ indicates lower is better. The result averaged over 10 seeds with each seed containing 300 episodes with $T_{\phi}=0.4$, $T_{min}=1$, $T_{max}=10$, and $\eta=0$
Mode Method Desired region $\uparrow$ Undesired region $\downarrow$ Score $\uparrow$ Preference Static $4.987\pm 0.649$ $0.284\pm 0.045$ Mixed Static $2.372\pm 0.123$ $000.000\pm 0$ 0.000 $0.739\pm 0.014$ Avoidance Static - $000.000\pm 0$ 0.000 $0.974\pm 0.011$
Table 13: Results correspond to the Highway environment. The up arrow $\uparrow$ indicates higher values are better and the down arrow $\downarrow$ indicates lower values are better. The hyper-parameters chosen are $T_{\phi}=0.6,T_{max}=5,T_{min}=0.3,\eta=0$
Mode $\alpha$ Desired lane $\uparrow$ Undesired lane $\downarrow$ Hits $\downarrow$ Score $\uparrow$ Avoidance 0.3 - $16.25\pm 3.45$ $0.15\pm 0.03$ $37.95\pm 0.68$ 0.5 - $5.80\pm 1.07$ $0.14\pm 0.23$ $37.44\pm 0.52$ 0.7 - $2.75\pm 0.49$ $0.47\pm 0.02$ $33.90\pm 0.91$ Preference 0.3 $9.64\pm 0.79$ $0.15\pm 0.02$ $37.58\pm 0.60$ 0.5 $11.58\pm 1.12$ $0.07\pm 0.06$ $38.14\pm 0.30$ 0.7 $23.74\pm 1.58$ $0.09\pm 0.01$ $38.42\pm 0.49$ Mixed 0.3 $11.90\pm 0.93$ $11.45\pm 1.27$ $0.13\pm 0.02$ $38.03\pm 0.55$ 0.5 $12.61\pm 1.20$ $12.36\pm 1.58$ $0.07\pm 0.01$ $36.97\pm 0.59$ 0.7 $23.18\pm 2.55$ $6.13\pm 1.47$ $0.11\pm 0.03$ $37.64\pm 0.51$